OpenStudy (anonymous):

2x^2+8+12=0 solve by using Quadratic formula

OpenStudy (anonymous):

recall that in general the quadratic equation is\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]which is used to solve equations of the given form\[ax^2+bx+c=0\]where a,b, and c are all constants, therefore you just plug in a=2, b=0, and c=20 into the quadratic equation and you will get your answer, just remember that\[\sqrt{-1}=i\]i being an imaginary number